what is 2 to the power of (12-2x)= 16 ? using logarithms
Answer by richwmiller(17219)

Well we know that 2^4=16
but we want to use logarithms.
so take the log in base 2 of both sides. Log 16 base 2 also happens to be 4
12-2x=4
8=2x
4=x
everything is coming up roses (4)
check
2^(12-2*4)=2^(12-8)=2^4=16
ok